Meeting Information

  • 1600 UT (1000 MDT)
  • AOC 317; IP:; phone: +1 575-835-7370


News, Travel, Visitors (McMullin)

  • Overall status: In A configuration (08 Jun); 4-band dipoles are removed tomorrow. 3-bit sampler testing is ongoing.
  • Key dates:
    • 07 May - 15 Jun: 4-band science observing
    • 30 Jun - Call for Proposals
    • 08 Jun - 12 Sep: A configuration (move from 31 May - 08 Jun).
    • 01 Aug - Proposal Deadline (C, CnB, B)
  • Staff Travel:
    • AM: 03 Jun - 18 Jun
    • LS: 9 May - 20 Jun
    • GM: 25 May - 24 Jun
    • FO: 12 Aug - 12 Sep
  • Visitors:
    • Robert Harris (AC982): 24 Apr - 16 Jul (High frequency w/ Claire)
    • Bob Sault: 16 May - 23 Jun (V Stokes; wobbles, splatter and other messy areas)
    • Michiel Brentjens: 11 Jun - 11 Nov (Holography; characterizing polarized beams)

Hardware status and updates (Langley, Perley, Rupen)

  • Array (Perley/Langley):
    • Antennas: EA21 back to array 11 Jul.
    • Receivers: EA12 X band; EA21 spate of receivers by the end of the month (28 Jun)
    • Calibration (Perley): Gain compression/expansion still under investigation - not understood. 3-bit sampler procurement has moved ahead; expect to have enough to populate 4 full antennas by August/September.
  • WIDAR (Rupen):
    • Notes (Rupen/Sowinski): 3-bit results summary; subarray testing.
  • Computing (Robnett):
    • New link to Lustre guidelines: (Robnett).
    • Notes: Lustre space is filling up; will need to handle past RSRO staff. Claire noted that more disks for existing desktops have been purchased - coming soon (20x2 TB disks). New nodes for the 2GC (2nd Generation Cluster) and cards for the Infiniband.

Software status and updates (Butler)

  • Links:
  • Sundry small updates (OST, TelCal); work toward preparation for D configuration. Demo on RCT for OSRO at 1pm today. Ongoing work with holography and added control on reference antenna and subbands.
  • CASA Update (new functionality available/issues/enhancements):Notes: New casapy-test (first 3.3 test version). Enhancements to opacity input for high frequency.
  • AIPS Update (new functionality available/issues/enhancements): Notes: Refinements to index table creation added.
  • SDM Update: Notes: Ephemeris table work is ongoing.

Commissioning Group Reports and Test Planning

  • NOTE: Please cc me on submitted tests so that it can be assigned the appropriate priority!

Commissioning priorities for the week (McMullin, Chandler)

Configuration move week (work around these activities:
  • Mon
    • 0900-1700: Testing/software development time
    • 1800-0900: Science operations and testing
  • Tue
    • 0900-1700: Testing/software development time
    • 1800-0900: Science operations and testing
  • Wed: Maintenance day
    • 0630-1700: Maintenance operations
    • 1800-0900: Science operations and testing
  • Thu:
    • 0900-1700: Testing/software development time
    • 1800-0900: Science operations and testing
  • Fri (not-yet-a-holiday):
    • 0900-1700: Testing/software development time
    • 1800-0900: Science operations and testing
  • Sat-Sun
    • Science operations + planned testing

Issues Affecting Science Observing:

  • Correlator: [Pokorny/Rupen/McMullin] time mislabeling in SDM data (non-synchronous operation between correlator dumps and CBE can require additional 1/2 integration time padding for online flags). This week!
  • Correlator: [Clark/Carlson] Delay rocking across bands. Implementation pending factor-of-2 reconciliation.
  • Correlator: [Rupen/Carlson w/ Momjian/Sault] V stokes issues.Data taken last night
  • Array: [Gustaaf] Amplitude wobbles seen on subset of antennas; phase jumps at scan boundaries also seen on subset of antennas (Ka, Q bands only thus far).
  • Array: [All] Zoo of other effects (some data are very good, some still seeing antenna-based IF/Polarization issue).

Testing This week/soon

  • CM/CBE update testing (subarrays+ deconfiguration problems).
  • 4-band testing (Owen): gasoline card reader interference issues. Concludes today.
  • phased array (Sowinski/McMullin): further tests from Amy/Ken (as above).Tests ready to roll
  • Array: *; Individual issue sleuthing (focus on phase jumps).
  • Astrometry (Bruthaler): UVW precision tests.
  • Correlator: Rupen: S-band shadow testing
  • Spectral line: Momjian: HI High resolution bandpass testing
  • V stokes testing (Sault/Momjian) TSPE0008_sb1850345); more tests coming

Commissioning Groups

  • Big deadline is end of this month for incorporation into links for Call for Proposals due 1 July.

    • Adv: Low band (Owen):
      • Notes:
    • Adv: Astrometry (Mioduszewski):
      • Notes: Begin testing of UVW precision issues (Brunthaler).
  • Calibrator Models
    • Notes: Status report and Ka/S update: 32 and 37 GHz observations available for 3C138, 3C147, 3C48 (3C286 by the end of the week); currently the information is put on the web (rather than added directly to CASA/AIPS; after A configuration data is added, it will be added directly. S band is not yet available.
  • HPC (Bhatnagar/Owen/Golap):
    • Notes:
  • Imaging (Ott/Golap):
    • Notes:
  • Algorithm (Bhatnagar, Owen):


RSRO A configuration status:

Individual reports (updates)


  • RSRO Calendar:
  • RSRO DnC, C config status:
  • Test Project Codes == *TRSR*
  • Guidelines for RSRO observing:
  • Notes:
    • Current Residents
      • D.J. Saikia (10C-119/AI145)::RSRO-75; 300G(2); x7106-POC: Momjian (Low Frequency)
      • J. Irwin (10C-119AI145)::RSRO-75; 300G(1); x7333 -- POC: Myers (Low Frequency)
      • Peter Hofner (10B-124/AH1011)::RSRO-26; 200G(1); x7496 -- POC: Ott (CASA testing)
      • Sergio Dzib (10B-131); 200G;x7496 -- POC: Mioduszewski (Astrometry)
      • Jacqueline van Gorkom (10C-211/AV324); 200D/x7098 -- POC: Momjian (Low Frequency)
      • Laura Chomiuk (*); 200B/x7497 -- POC: Goss (Bandpass stability)
      • Andreas Brunthaler::200E/x7246 -- POC: Mioduszewski (Astrometry/VLBI)
      • Jane Greaves::(AC982)::RSRO-4; 300G;x7106 -- POC: Chandler (High Frequency Documentation for Workshop)

Demo Science Status


  • EVLA Data Reduction Workshop 14-16 September 2011
Based on a report from the UC committee, it was decided that a workshop on EVLA data reduction should
be offered to support the community in transitioning from the early EVLA capabilities (consistent with the
historic VLA data volumes and rates) to the wider bandwidths/larger data volumes of the current RSRO
and future OSRO. I've attached the write-up on the history of the evolution of this workshop (below)
for those that are interested.

The workshop will be organized/coordinated by Gustaaf and the User Support group and is currently
slated for 14-16 September. The curriculum for the workshop will be developed based on our knowledge
of the current capabilities and is outlined below (only a half day on dedicated lectures, beyond that it
is hands on work with demo data):
- hardware/software issues
- data reduction techniques (currently available)
- special algorithms (particularly with regard to wide-band data)

I would like to hold a meeting to refine the curriculum and plan out the required training within our group
required to support this workshop. In addition, I would like to review what data sets are required to illustrate
the techniques (focused on D configuration processing, likely broken into high and low frequency groups).
  • 4-band summary:

4-band testing phases took place in:
   - 08 Sep - 29 Sep
        Six antennas deployed with dipoles:
                - EA03, EA11 (East)
                - EA01, EA19 (West)
                - EA06, EA28 (North)
        Problems with 3 (much smaller BWs seen); work to improve all; BWs measured used for OSS
        (only looking at 16 MHz).
   - 09 Feb - 23 Feb 4-band test installation
        Slow march to get all antennas working; by end of campaign had 60% working with the same
        response over 30 MHz, BW centered at 70 MHz. Work to get low-pass filters for all antennas
        to improve. Further tests in April prior to science campaign.
        Little interference seen (some narrow spikes at 70,78 MHz). 

Final mounting of the dipoles for the science campaign was 26 Apr - 15 Jun (for
science occuring from 07 May - 15 Jun).

   - 70/100 hours achieved; 11A-203, 11A-193 withdrawn 1st week of June based on problems outlined below.
Issues that compromised science goals of accepted programs:

   - Fundamentally, to reduce this data, the ionosphere must be removed which required detecting
        sources on the timescale of the atmosphere (roughly 30-60 seconds). This was not 
        achieved due to problems principally with unidentified broad band RFI:
   - RFI
        - birdies in band; this was able to be handled, however, some RFI mitigating strategies were
        deployed (shifting the sub-band to increase fringe rate).
        - broad-band RFI - this was the key problem that led to failures in the science goals.
        - tools were used to try to identify this (counter-rotate by the fringe rate of the
        correlator) to keep RFI phase constant and easier to pick out. However, sections of the
        bands would show RFI signals not readily correctable (see figure).
        This caused the marginal situation (calibration of the ionosphere) to break down and
        led to the inability to properly calibrate the data and achieve noise levels above
        what had been previously done.

Remaining tasks:
        - we would like to do an experiment with the card reader toggled off at the beginning
        and then on again in the middle to better quantify its impact. I've initiated discussions
        with Skip/Dan over this; they are currently planning to try to construct a cage around
        the reader to minimize its impact pending these results (apparently replacing it is not

Misc Actions

  • Old/Low priority
  • [McMullin]: Review/recover metrics for Priority 0 items.
  • [Perley/McMullin]: SPW/IF/BB lexicon and mapper
  • Development of WIDAR guidelines document (Michael)
-- JosephMcMullin - 2011-06-13
Topic revision: r4 - 2011-06-18, JosephMcMullin
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding NRAO Public Wiki? Send feedback